OpenStudy (anonymous):

what is 4 3/11 as a mixed number and in simplest form

OpenStudy (anonymous):

3/11 i think

OpenStudy (tkhunny):

Isn't that already a mixed number?

OpenStudy (anonymous):

yeah it is, he wrote the question wrong, but the answer should be 3/11

hero (hero):

A mixed number is a whole number plus a fraction. In other words a number of the form \(a \dfrac{b}{c} = a + \dfrac{b}{c}\) The other form it can be written is as an improper fraction. To write \(a \dfrac{b}{c} \) as an improper fraction, convert it to the form \(\dfrac{ac + b}{c}\)